$SPAD/src/input images6a.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994-1996.

———-

— * —

\getchunk{license}
-- 3-d surface whose height is the real part of the Gamma function,
-- and whose color is the argument of the Gamma function
gam(x,y) ==
  g := Gamma complex(x,y)
  point [x,y,max(min(real g, 4), -4), argument g]
draw(gam, -%pi..%pi, -%pi..%pi, title == "Gamma(x + %i*y)",
     var1Steps == 100, var2Steps == 100)
-- The Beta function
b(x,y) == Beta(x,y)
draw(b, -3.1..3, -3.1 .. 3, title == "Beta(x,y)")
-- The complex arc tangent function
atf(x,y) ==
  a := atan complex(x,y)
  point [x,y,real a, argument a]
draw(atf, -3.0..%pi, -3.0..%pi)

———-

References^{[1] nothing}

4